A.) What information is presented in the table above?

B.) Formulate the hypotheses for this test.

C.) Choose the appropriate forumla for the test statistic

D.) Describe the rejection region for this test

E.) What should the researcher conclude
Answer #1

a)

the table represent the frequency disrtibution of preference of sports of 40 randomely chosen Americans

b)

here null hypothesis: Ho: the data in sample fits the distribution claimed by newspaper article

Ha: the data in sample does not fits the distribution claimed by newspaper article

c)

applying chi square goodness of fit test:

observed Expected Chi square
category Probability(p) Oi Ei=total*p R2i=(Oi-Ei)2/Ei
Football 0.500 22.000 20.000 0.200
Baseball 0.300 14.000 12.000 0.333
Nascar 0.200 4.000 8.000 2.000
total 1.000 40 40 2.533

from above test statsitic \chi ^{2} =2.533

d)

for above degree of freedom =categories-1=3-1=2

for 2 df and 0.05 level rejection region  \chi ^{2}>5.991

e)

as test statitic is not in rejection region ; we can not reject null hypotheiss

we do not have sufficient evidence to reject article's claim.